# NoisyRollout: Reinforcing Visual Reasoning with Data Augmentation

### v1

@@ 3 Experiments
Table 1: Performance comparison of various VLMs on a suite of out-of-domain benchmarks. We report accuracy scores (%) for all benchmarks for clarity. “Avg.” denotes the average accuracy across the five benchmarks. Models marked with $\star$” are evaluated using our evaluation scripts implemented with vLLM (Kwon et al., 2023), and $\dagger$” indicates results evaluated using the direct-answer template. For models in the R1-related category, dataset sizes used for SFT and RL are annotated in red and blue, respectively. Among R1-related models, best value per column is bold, second best is underlined.
Model #Data MathVerse MathVision MathVista WeMath HallusionBench Avg.
Close-source
GPT-4o (Hurst et al., 2024) - 50.8 30.4 63.8 69.0 - -
Claude-3.5-Sonnet (Anthropic, 2024) - 26.5 38.0 67.7 - - -
Kimi1.5 (Kimi Team, 2025) - - 38.6 74.9 - - -
@@
InternVL-2.5-8B-Instruct (Chen et al., 2024) - 39.5 19.7 64.4 - - -
InternVL-2.5-78B-Instruct (Chen et al., 2024) - 51.7 32.2 72.3 - - -
QVQ-72B-Preview (Qwen, 2024) - - 35.9 71.4 - - -
Qwen2.5-VL-72B-Instruct (Bai et al., 2025) - - 38.1 74.8 - - -
URSA-8B (Luo et al., 2025) - 45.7 26.2 59.8 - - -
Mulberry-7B (Yao et al., 2024) - - - 63.1 - - -
R1-related (GRPO-tuned with rule-based reward)
R1-VL-7B (Zhang et al., 2025a) 260K+10K 40.0 24.7 63.5 - - -
Vision-R1-7B (Huang et al., 2025) 200K+10K 52.4 - 73.5 - - -
@@ 3.1 Main Results
Dataset & Evaluation. We use EasyR1 (Zheng et al., 2025) as our reinforcement learning training framework. Built on top of veRL (Sheng et al., 2024), EasyR1 is specifically designed to support VLMs. For our main experiments, we adopt the processed Geometry3K dataset (Lu et al., 2021) and the K12 dataset (Meng et al., 2025), each containing $2.1$K training samples. [^3] K12 originally contains approximately $8$K examples; we randomly sample $2.1$K to ensure consistency with Geometry3K in terms of training scale. To mitigate issues like reward hacking and model guessing, we convert all questions from multiple choice to free form formats. [^4] Frequently observed in our preliminary experiments, these issues might disrupt accurate reward assignment.
We evaluate model performance along two dimensions. First, we assess out-of-domain generalization across five benchmarks: four visual reasoning benchmarks, including MathVerse (Zhang et al., 2024a), MathVision (Wang et al., 2024), MathVista (Lu et al., 2023), and WeMath (Qiao et al., 2024), as well as one visual perception benchmark, HallusionBench (Guan et al., 2024). Second, we evaluate the in-domain performance of NoisyRollout by comparing it with the vanilla GRPO baseline on the Geometry3K test set. Moreover, we develop an evaluation suite for consistent assessment of our trained checkpoints and most open-source R1-related checkpoints using vLLM (Kwon et al., 2023) for accelerated inference (marked with $\star$), while adopting reported results for others. We employ greedy decoding with Gemini-2.0-Flash-001 (Gemini Team et al., 2023) as the judge to parse generated responses. While we carefully adopt the system prompts from respective codebases or papers, minor inconsistencies with reported results remain, which we deem acceptable given possible differences in judge models and evaluation implementations.
Implementation details. Following prior work (Yang et al., 2025; Huang et al., 2025; Deng et al., 2025), we initialize our policy model with Qwen2.5-VL-7B-Instruct, which exhibits strong foundational capabilities well-suited for subsequent RL training. In line with prior studies, we remove the KL divergence constraint with the reference model in GRPO to encourage exploration (Liu et al., 2025a), while keeping the vision encoder frozen for training stability and parameter efficiency (Meng et al., 2025). For other RL-related hyperparameters, we adopt the default settings from EasyR1, including a global batch size of $128$, a rollout batch size of $512$, a rollout temperature of $1.0$, and a learning rate of $61e{-}$ 6. Training is conducted over $15$ episodes and $60$ optimization steps ($T_{\mathrm{max}}$) on both datasets using $8$ A100-40G GPUs. For NoisyRollout-specific configurations, we implement Gaussian noise as our image distortion strategy. The initial Gaussian noise step $\alpha_{0}$ is set to $500$ for Geometry3K and $450$ for K12. We empirically set the rollout parameters to $n_{1}=n_{2}=6$, and apply a sigmoid-shaped noise annealing schedule:
$\alpha_{t}=\eta(\alpha_{0},t,T_{\text{max}})=\alpha_{0}\cdot\left(1-\frac{1}{1+e^{-\lambda(t/T_{\text{max}}-\gamma)}}\right)\textrm{,}$ (3)
where $\gamma=2/3$ controls the midpoint of the annealing curve, while $\lambda=30$ determines its steepness. This annealing schedule gradually reduces noise during training, helping to prevent convergence issues while preserving beneficial diversity in the early stages. Figure 6 visualizes the effect of applying different Gaussian noise steps to a clean image.
@@
Figure 3: Comparison of NoisyRollout and vanilla GRPO across in-domain and out-of-domain scenarios with the same total rollout number ($12$). The X-axis in all plots represents RL training steps. First column: Reward comparison on the in-domain dataset during training. Second and third columns: Comparison on four out-of-domain visual reasoning benchmarks. Last column: Evaluation of perception capabilities, where the upper subplot directly compares their perception performance on HallusionBench and the lower subplot presents the model-ranked Bradley–Terry win rates w.r.t. the perception qualities of their reasoning traces during training.
@@
Result 1: Out-of-domain generalization. When trained on the Geometry3K dataset, NoisyRollout not only improves in-domain performance (Figure 3, lower left subplot), but more importantly, demonstrates strong out-of-domain generalization. As shown in Table 1, NoisyRollout achieves $59.2\%$ average accuracy across five visual reasoning and perception benchmarks, consistently outperforming the GRPO baseline in every case. This advantage is further illustrated in Figure 3, which presents detailed comparisons across benchmarks as training progresses. Specifically, NoisyRollout achieves $53.2\%$ on MathVerse, $28.5\%$ on MathVision, and $69.6\%$ on WeMath, surpassing all R1-related models and even outperforming GPT-4o, which scores $50.8$% on MathVerse and $69.0$% on WeMath.
Moreover, NoisyRollout effectively mitigates perception degradation when reasoning templates are used: while Qwen2.5-7B-VL-Instruct’s perception accuracy drops from $71.2$% to $64.6$%, NoisyRollout not only avoids this degradation but improves performance to $72.1$%. The final subplot in Figure 3 confirms NoisyRollout delivers improved perception quality during reasoning, [^5] Gemini-2.0-Flash is used as the judge to assess perception quality, following carefully designed instructions and evaluation pipelines, based on 300 samples per benchmark. See Appendix A.3 for details. suggesting our hybrid rollout strategy enhances perceptual alignment despite training solely on clean images.
[…]
@@
Result 2: Sample efficiency. NoisyRollout demonstrates exceptional data efficiency by generalizing with only 2.1K training samples, whereas comparable models require significantly more data and additional SFT as warm-up training. For example, MM-Eureka-Qwen-7B uses 15K RL samples yet achieves only a $56.8\%$ average score, while R1-OneVision-7B needs 155K SFT samples and 10K RL samples but reaches just $52.0\%$. This efficiency stems from NoisyRollout’s use of noisy training signals that foster broader exploration during RL, enabling effective generalization from limited samples.
@@
Table 2: Performance comparison between NoisyRollout and vanilla GRPO on both out-of-domain and in-domain benchmarks under different rollout temperature settings. The total number of rollouts in all experiments is fixed at $12$. In vanilla GRPO, “$n(6):1.0$, $n(6):1.2$” indicates $6$ rollouts with temperature $1.0$ and another $6$ with temperature $1.2$. In NoisyRollout, “$n_{1}(6):1.0$” denotes $6$ rollouts per sample generated from clean input $(I,\mathbf{q})$ with temperature $1.0$, while “$n_{2}(6):1.0$” denotes $6$ rollouts per sample from noisy input $(\tilde{I},\mathbf{q})$ with temperature $1.0$. “Geo3K” represents the test set of Geometry3K dataset. The best value in each column is bold, and the second best is underlined.
@@
Method Temperature Geo3K MathVerse MathVision MathVista WeMath HallusionBench Avg.
@@
Result 3: Robustness across training datasets. NoisyRollout maintains strong generalization performance across diverse training sources. When trained on the K12 dataset, it achieves a $59.5\%$ average score, again surpassing the GRPO baseline ($57.9\%$). These results demonstrate NoisyRollout’s robustness across varying training sources and evaluation settings, confirming that its approach to reinforcement learning produces models with superior generalization capabilities for visual reasoning tasks.
Table 3: Performance comparison when using Dr. GRPO for optimization. “OOD Avg.” represents average accuracy (%) across five out-of-domain visual reasoning and perception benchmarks.
Method Geometry3K OOD Avg.
Qwen2.5-VL-7B-Instruct 39.4 53.3
+ Dr. GRPO 51.3 57.0
+ NoisyRollout (Dr. GRPO) 56.1 58.9
Result 4: Compatible with Dr. GRPO. Liu et al. (2025a) recently identified an optimization bias in GRPO that artificially increases response length. We evaluate NoisyRollout with their proposed Dr. GRPO’s unbiased objective and, as shown in Table 3, NoisyRollout maintains consistent improvements over the baseline, demonstrating that our hybrid rollout strategy’s benefits extend beyond the standard GRPO method.
@@ 3.2 Ablation Study: More Effective Rollout Diversity with Noisy Trajectories
Setup. We conduct ablation studies using Geometry3K as our training dataset (also applicable to the next subsection). In this part, we aim to examine the effectiveness of our NoisyRollout from the perspective of rollout diversity, a key factor for effective policy exploration in RL training. We define rollout diversity as the average pairwise cosine distance between trajectory embeddings, where higher values indicate greater diversity. We randomly sample 256 instances from the Geometry3K training set. For each sample, we generate either $n=12$ trajectories in vanilla GRPO or a combination of $n_{1}=6$ and $n_{2}=6$ trajectories in NoisyRollout, then encode them with an embedding model [^6] https://huggingface.co/sentence-transformers/all-MiniLM-L6-v2. We track both diversity and accuracy across training steps (Figure 4) and evaluate final performance on in-domain and out-of-domain benchmarks (Table 2).
Results. We use vanilla GRPO with a rollout temperature of $1.0$ as the control group. As shown in Figure 4, introducing noisy trajectories in NoisyRollout produces a similar effect to increasing the rollout temperature in vanilla GRPO (e.g., from $1.0$ to $1.2$): initial prediction accuracy on training samples decreases, but rollout diversity increases, ultimately leading to higher training accuracy in later stages. Moreover, both NoisyRollout and vanilla GRPO with a higher temperature ($1.2$) exhibit similar diversity dynamics compared to the control group — diversity is higher in the early stages of training, but drops below that of the control group later on. However, benchmark results in Table 2 show that NoisyRollout with temperature $1.0$ consistently outperforms vanilla GRPO across all tested temperatures ($0.8$ to $1.4$), as well as the mixed-temperature variant (using $6$ rollouts at temperature $1.0$ and $6$ at $1.2$). These results suggest that NoisyRollout introduces more effective rollout diversity from a visual perspective than simply adjusting temperature parameters, which increases diversity in a broad and less targeted way.
Table 4: Ablation study on total rollouts and noise steps. Left: Vanilla GRPO with total rollout $n$ versus our NoisyRollout with split rollouts $n_{1}$ and $n_{2}$. Right: Impact of initial noise steps on NoisyRollout performance. “OOD Avg.” reports average accuracy (%) across five out-of-domain visual reasoning and perception benchmarks. Best value per column is bold, second best is underlined.
Rollout Geometry3K OOD Avg.
$n=8$ 49.6 56.8
$n=12$ 51.4 57.2
$n=16$ 54.7 57.5
$n=20$ 53.6 57.3
$n_{1}=n_{2}=4$ 51.7 58.3
$n_{1}=n_{2}=6$ 54.9 59.2
$n_{1}=n_{2}=8$ 54.7 58.6
@@ 3.3 Ablation Study: Impact of Hyperparameters and Module Design
Noise annealing. As shown in Figure 5, removing noise annealing causes the in-domain performance of our method to drop sharply around training step $45$. This drop is due to divergence caused by a distributional mismatch—an issue discussed in Section 2 and further illustrated by the training dynamics in the same figure. Additionally, Table 5 shows that disabling noise annealing leads to lower performance in both in-domain and out-of-domain settings ($43.9\%$ and $58.0\%$, respectively), compared to our standard setting with noise annealing ($54.9\%$ and $59.2\%$). These results further highlight the effectiveness of noise annealing. We defer the discussion of other annealing strategies (e.g., power, exponential) to Appendix A.1.
Total rollout number. We analyze the impact of total rollout number by comparing vanilla GRPO and NoisyRollout under varying rollout budgets. As shown in Table 4 (left), increasing the number of rollouts in vanilla GRPO from $n=8$ to $n=16$ improves in-domain performance (from $49.6\%$ to $54.7\%$), but only marginally benefits out-of-domain generalization (from $56.8\%$ to $57.5\%$). In contrast, NoisyRollout consistently outperforms vanilla GRPO even when the total number of rollouts is held constant. Notably, NoisyRollout with $n_{1}=n_{2}=6$ (total $12$) achieves both higher in-domain ($54.9\%$) and out-of-domain ($59.2\%$) accuracy than vanilla GRPO with $16$ rollouts.
@@
Initial noise step. We evaluate the impact of noise strength by varying the initial Gaussian noise step, as shown in Table 4 (right). Gradually increasing the initial noise step $\alpha_{0}$ from 00 to $500$ consistently improves performance across all evaluation categories, suggesting that moderate noise promotes exploration and enriches the training signal. However, exceeding this threshold leads to performance degradation, as overly distorted images (see Figure 6) yield noisy rollouts with average near-zero rewards. These excessively noisy samples introduce harmful distribution shifts, ultimately destabilizing the learning process.
3.4 Unsuccessful Attempts
Noise image condition for both $\pi_{\theta_{\mathrm{old}}}$ and $\pi_{\theta}$. We attempted to optimize noisy rollouts based on their corresponding distorted images (rather than conditioning all policy updates on the clean image). This approach failed to show significant improvements over vanilla GRPO. We hypothesize that this strategy contradicts with the GRPO’s design, harming the information gain from the group advantage calculation, thus reducing their effectiveness for exploration during RL.
Image data augmentation. We explored a range of image distortions beyond Gaussian noise, including cropping and rotation. However, these augmentations often led to critical information loss, resulting in rollouts with consistently zero rewards. This caused unreliable policy gradient estimates and ultimately led to training instability and divergence. We also experimented with randomized augmentation, where one distortion (crop, rotation, or Gaussian noise) was randomly selected at each training step. However, this approach failed to improve stability. In contrast, Gaussian noise alone preserved essential visual information while introducing moderate perturbations. We find that such noise can act as a beneficial regularizer, injecting useful variability into the inputs.
Reward penalty on noisy subgroup. We experimented with applying explicit reward penalties (e.g., $-0.1$) to all noisy rollouts, aiming to encourage the model to better capture contrastive learning signals. However, this approach quickly led to training divergence. Rather than improving its core reasoning and perception abilities, the policy model learned to distinguish between clean and noisy rollouts. As a result, the noisy rollouts rapidly became highly off-policy, since the model could easily identify. This distributional mismatch destabilized training and undermined the learning.

### v3

@@ 3 Experiments
Dataset. We use EasyR1 [85] as our reinforcement learning training framework, which is built on verl [57] and specifically designed for VLMs. Our experiments utilize two datasets: Geometry3K [42], focused on geometric problem solving, and MMK12 [48], covering diverse K-12 math topics. These datasets comprise $2.1$K and $6.4$K training samples respectively. We processed them by converting all questions from multiple-choice to free-form format to prevent reward hacking and model guessing.
Evaluation. We mainly evaluate model performance along two dimensions. First, we assess out-of-domain generalization across five benchmarks: four visual reasoning benchmarks, including MathVerse [82], MathVision [64], MathVista [43], and WeMath [52], as well as one visual perception benchmark, HallusionBench [18]. Second, we evaluate the in-domain performance of NoisyRollout by comparing it with the vanilla GRPO baseline on the Geometry3K test set.
Moreover, we develop an evaluation suite for consistent assessment of our trained checkpoints and most open-source R1-related checkpoints using vLLM [28] for accelerated inference (marked with $\star$ in Tables 1 and 2), while adopting reported results for others. [^2] While we closely follow system (or format) prompts from relevant codebases or papers, minor result discrepancies may occur due to differences in judge models or inference engines, which we consider acceptable. We employ greedy decoding for model inference and use Gemini-2.0-Flash-001 [17] as the judge model to parse generated responses.
Table 1: Performance comparison of VLMs with moderate parameter sizes on a suite of out-of-domain benchmarks. Accuracy scores (%) are reported for all benchmarks for clarity. Models marked with $\star$” are evaluated using our evaluation suite. For R1-related models, the corresponding reasoning templates are used by default, while $\dagger$” indicates results obtained using the direct-answer template. Data sizes used for SFT and RL are annotated in blue and red, respectively. The best value in each column is shown in bold, and the second-best is underlined.
Model Data Size MathVerse MathVision MathVista WeMath HallusionBench
@@
InternVL-2.5-8B-Instruct [9] - 39.5 19.7 64.4 - 67.3$\dagger$
LLaVA-OneVision-7B [31] - 26.2 - 63.2 - 48.4$\dagger$
Kimi-VL-16B [27] - 44.9 21.4 68.7 - 66.2$\dagger$
URSA-8B [45] - 45.7 26.2 59.8 - -
Mulberry-7B [74] - - - 63.1 - -
R1-related (reinforcement learning with verifiable reward)
R1-VL-7B [81] 260K+10K 40.0 24.7 63.5 - -
Vision-R1-7B [21] 200K+10K 52.4 - 73.5 - -
@@
Table 2: Performance comparison of VLMs with large parameter sizes on a suite of out-of-domain benchmarks. The notation and evaluation protocols are consistent with those described in Table 1.
Model #Data MathVerse MathVision MathVista WeMath HallusionBench
Close-source
GPT-4o [22] - 50.8 30.4 63.8 69.0 71.4$\dagger$
Claude-3.5-Sonnet [3] - 26.5 38.0 67.7 - 71.6$\dagger$
Kimi1.5 [26] - - 38.6 74.9 - -
Open-source
InternVL-2.5-78B-Instruct [9] - 51.7 32.2 72.3 - 72.9$\dagger$
QVQ-72B-Preview [53] - - 35.9 71.4 - -
Qwen2.5-VL-72B-Instruct [4] - - 38.1 74.8 - 71.9$\dagger$
R1-related (RL-tuned with verifiable reward)
MM-Eureka-Zero-38B [48] 9.4K 48.9 26.6 64.2 - -
[…]
Implementation details. Following prior work [48, 67], we initialize our policy models with Qwen2.5-VL-7/32B-Instruct, which exhibit strong foundational capabilities well-suited for subsequent RL training. All experiments are conducted using $8$ A100 GPUs (40G for $7$B model, 80G for $32$B model). We keep the vision encoder frozen for training stability and parameter efficiency. For other general RL-related hyperparameters, we adopt the default settings from EasyR1: a global batch size of $128$, a rollout batch size of $512$, a rollout temperature of $1.0$, and a learning rate of $1\text{e}{-}6$. To prevent token-length bias, we compute the policy loss using the token-mean aggregation strategy. [^3] The code implementation can be found at verl. For NoisyRollout-specific configurations, we adopt Gaussian noise as the image distortion strategy, and apply a sigmoid-shaped annealing schedule:
$\alpha_{t}=\eta(\alpha_{0},t,t_{\text{max}})=\alpha_{0}\cdot\left(1-\frac{1}{1+e^{-\lambda(t-\gamma)/t_{\text{max}}}}\right),$ (3)
where $\gamma$ determines the midpoint of the annealing curve and $\lambda$ controls its steepness. Figure 7 illustrates the visual effects of applying different levels of Gaussian noise to a clean image. We defer the discussion of image distortion strategies (e.g., cropping, rotation), noise annealing strategies (e.g., power, exponential), and proportions of noisy rollouts in total rollouts to Appendix A. The reasoning and direct-answer templates used in our experiments are shown in Appendix I. Additional implementation details regarding the number of training steps/epochs and the hyperparameters for image distortion and noise annealing are presented in Appendix K.
@@ 3.1 Main Results
Figure 3: Comparison of NoisyRollout and vanilla GRPO on Qwen2.5-VL-7B-Instruct across in-domain and out-of-domain scenarios with the same total rollout number ($12$). The X-axis in all subplots represents RL training steps. First column: Reward comparison on the in-domain dataset during training. Second and third columns: Comparison on four out-of-domain visual reasoning benchmarks. Last column: Evaluation of visual perception capabilities, where the upper subplot directly compares their perception performance on HallusionBench and the lower subplot presents the model-ranked Bradley–Terry win rates w.r.t. the perception qualities of their reasoning traces.
@@
Result 1: Out-of-domain generalization. When trained on the Geometry3K dataset using Qwen2.5-VL-7B-Instruct, NoisyRollout not only improves in-domain performance (Figure 3, lower left subplot), but more importantly, demonstrates strong out-of-domain generalization. As shown in Table 1, NoisyRollout achieves superior performance across five visual reasoning and perception benchmarks, consistently outperforming the vanilla GRPO baseline in every case. This advantage is further illustrated in Figure 3, which presents detailed comparisons across benchmarks as training progresses. Specifically, NoisyRollout achieves $53.2\%$ on MathVerse, $28.5\%$ on MathVision, and $69.6\%$ on WeMath, surpassing existing R1-related baselines and even outperforming GPT-4o.
Moreover, while Qwen2.5-7B-VL-Instruct’s perception accuracy on HallusionBench drops from $71.2$% to $64.6$% when switching from direct-answer to reasoning templates, [^4] This degradation caused by the reasoning template has also been observed in previous studies [6, 25]. NoisyRollout achieves $72.1$% with the reasoning prompt (compared to vanilla GRPO’s $69.8$%). The final subplot in Figure 3 further confirms that NoisyRollout enhances perception quality during reasoning, achieving a higher Bradley–Terry win rate over vanilla GRPO (See Appendix D for details). These results indicate that our hybrid rollout strategy enhances visual perception by promoting better policy exploration through vision-oriented inductive biases.
@@
Result 2: Sample efficiency. NoisyRollout demonstrates exceptional data efficiency by generalizing with only 2.1K training samples from Geometry3K, whereas comparable models require significantly more data or even additional SFT as warm-up training. For example, Table 1 indicates that OpenVLThinker-7B needs 35K SFT samples and 15K RL samples but reaches only $48.0\%$ on MathVerse and $71.5$% on MathVista. This efficiency stems from NoisyRollout’s use of noisy training signals that foster targeted exploration during RL, enabling effective generalization from limited samples.
Result 3: Robustness across training datasets and model sizes. NoisyRollout consistently improves upon vanilla GRPO, demonstrating strong robustness across model sizes and training datasets. As shown in Tables 1 and 2, the $7$B model trained on MMK12 achieves gains of $1.2$%, $1.3$%, and $1.9$% over GRPO on MathVerse, MathVista, and HallusionBench, respectively. Similarly, the 32B model trained on MMK12 surpasses GRPO by $0.4$%, $0.7$%, and $1.1$% on the same benchmarks.
Notably, in certain benchmarks like MathVerse and MathVista, the performance gains of NoisyRollout over vanilla GRPO are smaller for the $32$B model than for the $7$B model. This is likely because the $32$B model’s initial policy was already fine-tuned via RL, [^5] https://qwenlm.github.io/blog/qwen2.5-vl-32b/ whereas the 7B model’s was not.
@@
Table 3: Performance comparison under different rollout temperature settings, with the total number of rollouts fixed at $12$. In vanilla GRPO, “$n(6):1.0$, $n(6):1.2$” indicates $6$ rollouts with temperature $1.0$ and another $6$ with temperature $1.2$. In NoisyRollout, “$n_{1}(6):1.0$” denotes $6$ rollouts per sample generated from clean input $(I,\mathbf{q})$ with temperature $1.0$, while “$n_{2}(6):1.0$” denotes $6$ rollouts per sample from noisy input $(\tilde{I},\mathbf{q})$ with temperature $1.0$. “Geo3K” represents the test set of Geometry3K dataset. “Avg.” represents average accuracy (%) across six benchmarks.
@@
Method Rollout Temperature Geo3K MathVerse MathVision MathVista WeMath HallusionBench Avg.
@@ 3.2 Ablation Study: More Effective Rollout Diversity with Noisy Trajectories
Setup. Unless otherwise specified, all ablation studies in this and the following subsection use Geometry3K as the training dataset on Qwen2.5-VL-7B-Instruct. In this part, we aim to examine the effectiveness of our NoisyRollout from the perspective of rollout diversity, a key factor for effective policy exploration in RL training. Here, we define rollout diversity as the average pairwise cosine distance between trajectory embeddings, where higher values indicate greater diversity. We randomly sample 256 instances from the Geometry3K training set. For each sample, we generate either $n=12$ trajectories in vanilla GRPO or a combination of $n_{1}=6$ and $n_{2}=6$ trajectories in NoisyRollout, then encode them with an embedding model. [^6] https://huggingface.co/sentence-transformers/all-MiniLM-L6-v2 We track both diversity and accuracy across training steps (Figure 4) and evaluate final performance on in-domain and out-of-domain benchmarks (Table 3). We use vanilla GRPO with a rollout temperature of $1.0$ as the control group.
Result. As shown in Figure 4, NoisyRollout enhances rollout diversity in early training stages compared to the control group, similar to increasing rollout temperature in vanilla GRPO from $1.0$ to $1.2$. This initial diversity boost, though accompanied by lower starting accuracy, ultimately leads to higher final training accuracy. Moreover, both NoisyRollout and higher-temperature vanilla GRPO show diversity decreasing below the control group in later training stages.
Table 3 reveals that NoisyRollout with temperature $1.0$ consistently outperforms vanilla GRPO across all temperature settings ($0.8$ to $1.4$), as well as mixed-temperature variants. Moreover, when applying temperature $1.2$ to both approaches, NoisyRollout still demonstrates significant improvement over vanilla GRPO. These results indicate that NoisyRollout introduces more targeted and effective diversity than simply adjusting temperature parameters, which increases diversity in a less focused manner.
Figure 6: Performance comparison on MMK12 when scaling up the training data size.
Table 4: Ablation study on the impact of initial noise steps ($\alpha_{0}$). “OOD Avg.” represents average accuracy (%) across five out-of-domain benchmarks.
@@ 3.3 Ablation Study: Impact of Hyperparameters and Module Design
Data scale. Although Geometry3K is a high-quality training dataset, its limited size ($2.1$K samples) prevents a thorough investigation of the scaling behavior of NoisyRollout compared to vanilla GRPO. To enable such analysis, we additionally consider MMK12, which contains $6.4$K samples after preprocessing. Figure 6 shows NoisyRollout consistently outperforms vanilla GRPO across various data scales, ranging from $1.1$K to $6.4$K. Notably, the performance gains do not diminish as the dataset size increases, suggesting that NoisyRollout has strong potential for use in large-scale training regimes.
@@
Noise annealing. As shown in Figure 5, removing noise annealing causes the in-domain performance of our method to drop sharply around training step $45$. This drop is due to divergence caused by a distributional mismatch—an issue discussed in Section 2 and further illustrated by the training dynamics in the same figure. Additionally, Table 5 shows that disabling noise annealing leads to lower performance in both in-domain and out-of-domain settings ($43.9\%$ and $58.0\%$, respectively), compared to our standard setting with noise annealing ($54.9\%$ and $59.2\%$). These results further highlight the effectiveness of noise annealing.
@@
Initial noise step. We evaluate the impact of noise strength by varying the initial Gaussian noise step, as shown in Table 4. Gradually increasing the initial noise step $\alpha_{0}$ from 00 to $500$ consistently improves performance across all evaluation categories, suggesting that moderate noise promotes exploration and enriches the training signal. However, exceeding this threshold leads to performance degradation, as overly distorted images (see Figure 7) yield noisy rollouts with average near-zero rewards. These excessively noisy samples introduce harmful distribution shifts during policy updates, ultimately destabilizing the learning process. Additional ablation results on the MMK12 dataset are deferred to Appendix A. [^7] We also include additional ablations (e.g., number of rollouts and data seed variations) in Appendix A.
Figure 8: Comparison of gradient projection ratio over training.
Table 6: Performance comparison when using GRPO variants for policy optimization.
Method Geometry3K OOD Avg.
Qwen2.5-VL-7B-Instruct 39.4 53.3
+ GRPO (w.o. $\mathrm{std}(\mathbf{r})$) 51.3 57.0
+ NoisyRollout (w.o. $\mathrm{std}(\mathbf{r})$) 56.1 58.9
+ GRPO ($\epsilon_{\mathrm{high}}=0.28$) 52.6 58.2
+ NoisyRollout ($\epsilon_{\mathrm{high}}=0.28$) 53.9 59.6
GRPO variant. Recently, several variants have been proposed to enhance the original GRPO implementation. Specifically, Liu et al. [40] identified a question-level difficulty bias and proposed removing the standard deviation normalization ($\mathrm{std}(\mathbf{r})$) to address this issue. In addition, Yu et al. [76] increased the upper clipping threshold ($\epsilon_{\mathrm{high}}$) to mitigate entropy collapse. As shown in Table 6, applying NoisyRollout consistently improves performance not only on the original GRPO implementation but also across these variants. This highlights that NoisyRollout provides complementary benefits alongside optimization-focused modifications, underscoring its broad applicability.
3.4 Further Analysis: Quantitative Contribution of Noisy Rollouts on RL Optimization
Setup. For each training sample, we partition the collected rollouts into Clean and Noisy subgroups, containing $n_{1}$ and $n_{2}$ rollouts, respectively. We measure each subgroup’s contribution by projecting its specific effective gradients onto an anchor gradient $\mathbf{g}^{t}=\theta^{t+\Delta t}-\theta^{t}$, which represents the overall model update over $\Delta t$ optimization steps, beginning at training step $t$. The subgroup effective gradients, $\mathbf{g}^{t}_{\text{clean}}$ and $\mathbf{g}^{t}_{\text{noisy}}$, are derived from actual optimization steps starting from $\theta^{t}$, using only rollouts from the respective subgroup (by masking losses from the other subgroup). The projection ratios are then calculated as $r^{t}_{\text{clean}}=(\mathbf{g}^{t}_{\text{clean}}\cdot\mathbf{g}^{t})/\|\mathbf{g}^{t}\|^{2}$ and $r^{t}_{\text{noisy}}=(\mathbf{g}^{t}_{\text{noisy}}\cdot\mathbf{g}^{t})/\|\mathbf{g}^{t}\|^{2}$. These ratios provide a quantitative estimate of each subgroup’s contribution to the overall model update $\mathbf{g}^{t}$. More details are included in Appendix J.
Result. Figure 8 shows that the Noisy subgroup consistently contributes more significantly to policy optimization compared to Clean, especially during early training phases when distortion strength $\alpha_{t}$ is high and the policy $\pi_{\theta}$ still struggles with visual understanding. This trend gradually diminishes towards the final stages of training as the learning is gradually “on-policy”. These findings quantitatively confirm that our method effectively leverages noisy rollouts to enhance training signals.